# On a conjecture of Peter Neumann on fixed points in permutation groups

### v1

@@ 8.7. Class $\mathcal{C}_{7}$
Assume now $G=\mathrm{Sp}_{n}(q)$ and $M$ of type $\mathrm{Sp}_{m}(q)\wr S_{t}$, with $qt$ odd. Then we may choose an element as above but where $x$ has order $q^{m/2}+1$. Then, the dimension of the centralizer of $g$ in $\mathrm{Sp}_{n}(K)$ is at most $m^{2t-2}/2+n/2$ (see [56, Theorem 3.1]). Moreover, $g$ has $n/m=m^{t-1}$ generalized Jordan blocks over $K$, so by Lemma 7.11(iii) we have $|C_{G}(g)|<2^{m^{t-1}}q^{m^{2t-2}/2+n/2}\leq q^{m^{t-1}+m^{2t-2}/2+n/2}$, and so
@@
which is less than $|G|$ if $n>8$ (i.e., $n\geq 32$). If $n=8$, since $M\in A$ we have $q\geq 5$ (see [7]), and $M=\mathrm{Sp}_{2}(q)^{\otimes 3}.2^{2}.S_{3}$ and $g$ has two $1$-Jordan blocks of dimension $2$, and one rational $f$-block of dimension $4$, where $f$ is the minimum polynomial of an element of order $(q+1)/2$ over $\mathbb{F}_{q}$. (A rational $f$-block refers to an indecomposable $\mathbb{F}_{q}\langle g\rangle$-submodule whose minimum polynomial is $f$.) In particular, the dimension of the centralizer of $g$ in $\mathrm{Sp}_{n}(K)$ is $2+4=6$ (see [56, Theorem 3.1]), and $|C_{G}(g)|=q^{4}|\mathrm{GU}_{1}(q)||\mathrm{O}^{\pm}_{2}(q)|\leq 2q^{4}(q+1)^{2}$ and we see that $|M||C_{G}(g)|^{3}<|G|$.
The other cases are similar. These are $G=\mathrm{SU}_{n}(q)$ and $M$ of type $\mathrm{GU}_{m}(q)\wr S_{t}$; $G=\Omega^{+}_{n}(q)$ and $M$ of type $\mathrm{O}_{m}^{\pm}(q)\wr S_{t}$ ($q$ odd) or $\mathrm{Sp}_{m}(q)\wr S_{t}$ (qtqt even); $G=\Omega_{n}(q)$ with $qn$ odd and $M$ of type $\mathrm{O}_{m}(q)\wr S_{t}$. Except for the case where $M$ is of type $\mathrm{O}_{m}^{\pm}(q)\wr S_{t}$, we can choose $g=x\otimes\cdots\otimes x\otimes y$ where $x$ is semisimple with distinct eigenvalues and $y$ unipotent with one Jordan block, as above. In the exceptional case, we have $m\geq 4$ and we choose $y$ unipotent with two Jordan blocks of size $m/2$. We have that all generalized Jordan blocks of $g$ over $K$ have size $m/2$, and each such block occurs with multiplicity at most $2m^{t-2}$, so the dimension of the centralizer of $g$ in $\mathrm{GL}_{n}(K)$ (resp. $\mathrm{O}_{n}(K)$) is at most $2m^{2t-2}$ (resp. at most $m^{2t-2}$, see [56, Theorem 3.1]), and the conclusion follows. ∎
8.8. Class $\mathcal{C}_{8}$
@@
The cases $G=\mathrm{SL}_{n}(q)$ and $M$ of type $\mathrm{O}^{-}_{n}(q)$ or $\mathrm{Sp}_{n}(q)$ ar similar; in both cases we choose $g$ of order $q^{n/2}+1$. Finally, for $G=\mathrm{Sp}_{n}(q)$ and $M=\mathrm{O}^{\varepsilon}_{n}(q)$ with $\varepsilon\in\{+,-\}$ and $q$ even, letting $g\in M$ be an element without eigenvalue $1$, we have $\mathrm{fix}(g,G/M)=1$. ∎
@@ 8.9. Class $\mathcal{N}$
Assume first $G=\mathrm{P}\Omega^{+}_{8}(q)$; we go through the possibilities given in [7, Table 8.50], which is taken from [36]. For convenience, we work in $G=\Omega^{+}_{8}(q)$, and put d=(2,q1)d=(2,q-1). Using the notation as in the table, we are only concerned with novelties $\mathrm{N_{1}}$ and $\mathrm{N_{4}}$. If $M$ is parabolic, we use Lemma 6.3. Assume now $M\cong d\times G_{2}(q)$. We choose $g\in\mathrm{SL}_{3}(q)<M$ of order a ppd of $q^{3}-1$. Then $g$ preserves a nondegenerate $6$-space $W$ of plus type (acting irreducibly on complementary maximal totally singular subspaces), so $|C_{G}(g)|\leq(q^{3}-1)(q-1)$. Moreover $|C_{M}(g)|=d(q^{3}-1)/(q-1)$ and $g^{G}\cap M=g^{M}$ (note all cyclic subgroups of $G_{2}(q)$ of order $|g|$ are conjugate). Therefore
@@
Assume next $M=(\Omega^{+}_{2}(q)\times\frac{1}{d}\mathrm{GL}_{3}(q)).[2d]$ with $q\geq 3$; then $M$ is the stabilizer of $\{A,B\}$ where $A$ and $B$ are totally singular $3$-spaces with trivial intersection and $A\oplus B$ nondegenerate; see [36, proof of Proposition 3.2.3]. The subgroup $H=\Omega^{+}_{2}(q)\times\frac{1}{d}\mathrm{GL}_{3}(q)$ preserves $A$ and $B$. We let $g\in\mathrm{SL}_{3}(q)<H$ be of order a ppd of $q^{3}-1$, acting trivially on $(A\oplus B)^{\perp}$. Note that if $g$ lies in a conjugate of $M$, then it lies in the corresponding conjugate of $H$; but $g$ fixes precisely two totally singular $3$-spaces and so $\mathrm{fix}(g,G/M)=1$.
Next let $M=(\Omega^{-}_{2}(q)\times\frac{1}{d}\mathrm{GU}_{3}(q)).[2d]$ with $q\geq 3$. The subgroup $H=\Omega^{-}_{2}(q)\times\frac{1}{d}\mathrm{GU}_{3}(q)$ is the stabilizer in $\mathrm{GU}_{4}(q^{2})\cap G$ of a nondegenerate $1$-space for the unitary geometry (see [36, proof of Proposition 3.2.2]). We let $g\in H$ be of order a ppd of $q^{6}-1$ if $q\neq 2$, and of order $9$ if $q=2$. Similarly to the previous case we have $\mathrm{fix}(g,G/M)=1$.
Assume finally $M=(D_{2(q^{2}+1)/d})^{2}[2d].S_{2}$. Then $M$ is the normalizer of a Sylow $r$-subgroup of $G$, where $r$ is an odd prime divisor of $q^{2}+1$; we have that $M$ stabilizes a decomposition $V=V_{1}\perp V_{2}$ into $4$-spaces of minus type (see [36, Proof of Proposition 3.3.1]). Let $g\in M$ be of order $(q^{2}+1)/d$, acting irreducibly on $V_{1}$ and centralizing $V_{2}$. Then $\mathrm{fix}(g,G/M)$ is equal to the number of Sylow $r$-subgroups of $\Omega(V_{2})\cong\mathrm{PSL}_{2}(q^{2})$, which is $q^{2}(q^{2}-1)/2$, and
@@
Assume now $G=\mathrm{Sp}_{4}(q)$ with $q$ even; the classes in $\mathcal{N}$ can be found in [2, Section 14] (see also [7]). If $M$ is parabolic, then we use Lemma 6.3. If $M=(C_{q-1})^{2}\colon D_{8}$, then $M$ is the normalizer of a nondegenerate $2$-space of plus type in $\mathrm{SO}^{+}_{4}(q)<G$. We let $g\in C_{q-1}^{2}<M$ be of order q1q-1 with eigenvalues $\lambda^{\pm 1},\lambda^{\pm 2}$ on the natural module $V$; then $\mathrm{fix}(g,G/M)=1$, since $g$ fixes only two nondegenerate $2$-spaces. Assume next $M=(C_{q+1})^{2}\colon D_{8}$, so $M$ is the normalizer of a nondegenerate $2$-space of minus type in $\mathrm{SO}^{+}_{4}(q)<G$. We let $g\in M$ be of order $q+1$, with distinct eigenvalues on $V$, and similarly to the previous case we have $\mathrm{fix}(g,G/M)=1$. Assume finally $M=C_{q^{2}+1}\colon 4$, so $M$ is the normalizer of a subgroup $C_{q^{2}+1}$. Then we let $g\in M$ be of order $q^{2}+1$, and we have $\mathrm{fix}(g,G/M)=1$ as in the proof of Lemma 8.6. The proof is now complete. ∎
@@ 9. Classical groups: class $\mathcal{S}$
In this section we complete the proof of Theorem 3.1. The only remaining case is when $G$ is classical and $M$ is in the class $\mathcal{S}$ of subgroups. Throughout this section, $V=\mathbb{F}_{q}^{n}$ denotes the natural module for $G$, where $q$ is a power of the prime $p$. (This is a change of notation from Section 8, where the natural module is $\mathbb{F}_{q^{u}}^{n}$ – so in this section, the unitary case is $G=\mathrm{PSU}_{n}(q^{1/2})$.) Class $\mathcal{S}$ consist of the members $M$ of $\mathcal{A}(G)$ that do not belong to $\mathcal{C}_{1},\ldots,\mathcal{C}_{8},\mathcal{N}$. It was proved by Aschbacher [2] (see also [49]) that these subgroups $M$ are almost simple. Moreover, if $L$ denotes the quasisimple cover of $S=\text{Soc}(M)$ acting faithfully on $V$, then $V$ is an absolutely irreducible $\mathbb{F}_{q}L$-module, which cannot be realized over a proper subfield of $\mathbb{F}_{q}$, and finally, if $G=\mathrm{PSL}_{n}(q)$ then $L$ fixes no nondegenerate unitary or bilinear form on $V$.
@@ 9.2. Some preliminary lemmas
Let $a$ be the dimension of the $1$-eigenspace of $g$, let $d$ be the dimension of the centralizer of $g$ in $\mathrm{GL}_{n}(K)$ and let $E_{2}$ be the number of distinct irreducible factors of degree at least $2$ of the characteristic polynomial of $g$. By Lemmas 7.10 and 7.11 we have $|C_{G}(g)|\leq q^{d/2+a/2+E_{2}}$. Next, note that $E_{2}$ is at most half the number of nontrivial distinct eigenvalues of $g$. Assume first (9.5), so by Lemma 7.7 (with $A=cn$) we have $d+a+2E_{2}\leq B+n^{2}/B$. Therefore, by Lemma 7.9 we get
@@ 9.3. The proof
We now fix some notation. Recall that $G$ is a simple classical group with natural module $V=\mathbb{F}_{q}^{n}$, where $q$ a power of the prime $p$. We set $K=\overline{\mathbb{F}_{p}}$. We have that $M\in\mathcal{A}$ is in class $\mathcal{S}$, so $M$ is almost simple; we denote by $S$ the socle of $M$. We will find it convenient, in the proofs, to replace $G$ by the quasisimple cover acting faithfully on $V$, namely $\mathrm{SL}_{n}(q)$, $\mathrm{SU}_{n}(q^{1/2})$ or $\mathrm{Sp}_{n}(q)$ or $\Omega^{\varepsilon}_{n}(q)$. We will still denote by $M$ the preimage in $\mathrm{SL}_{n}(q)$, while we will denote by $L\leq\mathrm{SL}_{n}(q)$ a quasisimple cover of $S$ with $S=L/Z(L)$. For our choice of the element $g\in L$, we will keep throughout the following notation:
@@
Note that $M$ preserves a nondegenerate quadratic or bilinear form on $V$. Since $M\in\mathcal{A}$, it follows that $G$ is orthogonal or symplectic over $\mathbb{F}_{p}$; see for example [39, p. 187] for the precise embeddings. We ignore the cases $G=\mathrm{Sp}_{4}(2)^{\prime}\cong A_{6}$, $G=\Omega^{-}_{4}(2)\cong\mathrm{SL}_{2}(4)$, $G=\Omega_{5}(2)\cong\mathrm{PSp}_{4}(2)$, andqw $G=\Omega^{\pm}_{6}(2)\cong\mathrm{PSL}^{\pm}_{4}(2)$, which have been handled in Lemma 8.1. Assume first $n=4$, so by [7] we have $G=\Omega^{-}_{4}(q)\cong\mathrm{PSL}_{2}(q^{2})$ and $M=A_{5}$ with $q=p\geq 7$. Let $g\in M$ be of order $3$, so $|C_{G}(g)|=(q^{2}-1)/2$ and by Lemma 2.1 we have $\mathrm{fix}(g,G/M)=(q^{2}-1)/6<|G:M|^{1/3}$. Assume now $n=5$, so $G=\Omega_{5}(q)\cong\mathrm{PSp}_{4}(q)$ and either $A_{6}\leq M\leq S_{6}$ with $q=p\geq 5$ or $M=A_{7}$ with $q=p=7$. If $M=A_{7}$ we let $g\in M$ be of order $7$, so $g$ is regular unipotent and $|C_{G}(g)|=q^{2}$ and so $\mathrm{fix}(g,G/M)\leq 2q^{2}/7<|G:M|^{1/3}$. If $M\leq S_{6}$ then we let $g\in M$ be of order $5$, so $g$ is regular in $\mathrm{SL}_{n}(q)$ and $|C_{G}(g)|\leq(q+1)^{2}$ and $\mathrm{fix}(g,G/M)\leq 2(q+1)^{2}/5<|G:M|^{1/3}$. Assume then $n=6$, so $G=\Omega^{\pm}_{6}(q)\cong e.\mathrm{PSL}^{\pm}_{4}(q)$ and $M=e\times A_{7}$ with $q=p\geq 11$ (here $e\in\{1,2\}$). Let $g\in M$ be of order $7$, so $|C_{G}(g)|\leq e(q+1)^{3}/2$; moreover $g^{G}$ splits into at most two $M$-classes and so by Lemma 2.1 we have $\mathrm{fix}(g,G/M)\leq(q+1)^{3}/7<|G:M|^{1/3}$. The case $n\geq 7$ is analogous. Specifically, let $g\in A_{m}$ be an $(m-\delta)$-cycle, where $\delta=1$ if $m$ is even and $\delta=0$ if $m$ is odd. In both cases, $g$ is regular in $\mathrm{SL}_{n}(q)$, and $g^{G}$ splits into at most two $M$-classes, so $\mathrm{fix}(g,G/M)\leq 2|C_{G}(g)|/(m-\delta)$, which is easily seen to be $|G|^{1/3}/((2,q-1)\cdot m!)^{1/3}\leq|G:M|^{1/3}$. For example, for $n=8$ and $G=\mathrm{Sp}_{8}(2)$ and $A_{10}\leq M$, $g$ acts on $\mathbb{F}_{2}^{8}$ with irreducible submodules of dimensions $2$ and $6$, so $|C_{G}(g)|=(2^{3}+1)(2+1)=27$ and
@@
If $n$ satisfies the above inequalities, then by Lemmas 9.1, 9.2, 9.3 and 9.4 we have $|C_{G}(g)|<|G:M|^{1/3}$, as desired. Assume then that $n$ is smaller than the above values. Then, by [32], we see that $m\leq 14$. If $m\leq 13$, then the modular character tables of $L$ are available in [8] and the result follows. If $m=14$, then the modular character tables are available unless $L=2.A_{14}$. By [32], we see that in this case the only option is $n=32$, $p=7$ and $G=\mathrm{Sp}_{32}(q)$. By Lemma 7.17 we have that $L=A_{14}$ is generated by two conjugates of an element $g$ of order $11$. We deduce that either $a\leq 12$ and each nontrivial eigenspace has dimension at most $2$, or $a\leq 2$ and each nontrivial eigenspace has dimension at most $3$; it follows that $d\leq 184$, which implies $|C_{G}(g)|<|G:M|^{1/3}$. ∎
@@
Assume first $S$ is in $\mathcal{H}_{1}$. Suppose $S$ is in $\mathcal{H}_{1}$(1); then $L=S$. If $S$ is sporadic, let $g=x_{1}$ be as in Table 7.2. If $S$ is of Lie type and $S\neq\mathrm{PSL}^{\varepsilon}_{m}(r)$, let $g=x_{1}$ be as in Table 7.1. If $S=\mathrm{PSL}^{\varepsilon}_{m}(r)$, set $u=1$ if $\varepsilon=+$ and $u=2$ if $\varepsilon=$-, and let $g$ be an element of order $\Phi^{*}_{um}(r)$ if $m$ is odd, and of order $\Phi^{*}_{u(m-1)}(r)$ if $m$ is even. Then, in all cases, [27, Theorem 1.1, Props 3.4–3.13, Prop 4.5] shows that $L$ is generated by three conjugates of $g$ whose product is $1$. It follows from Lemma 7.3 that $a\leq n/3$, as desired.

### v2

@@ 8.7. Class $\mathcal{C}_{7}$
Assume now $G=\mathrm{Sp}_{n}(q)$ and $M$ of type $\mathrm{Sp}_{m}(q)\wr S_{t}$, with qtqt odd. Then we may choose an element as above but where $x$ has order $q^{m/2}+1$. Then, the dimension of the centralizer of $g$ in $\mathrm{Sp}_{n}(K)$ is at most $m^{2t-2}/2+n/2$ (see [56, Theorem 3.1]). Moreover, $g$ has $n/m=m^{t-1}$ generalized Jordan blocks over $K$, so by Lemma 7.11(iii) we have $|C_{G}(g)|<2^{m^{t-1}}q^{m^{2t-2}/2+n/2}\leq q^{m^{t-1}+m^{2t-2}/2+n/2}$, and so
@@
which is less than $|G|$ if $n>8$ (i.e., $n\geq 32$). If $n=8$, since $M\in A$ we have $q\geq 5$ (see [7]), and $M=\mathrm{Sp}_{2}(q)^{\otimes 3}.2^{2}.S_{3}$ and $g$ has two 11-Jordan blocks of dimension 22, and one rational $f$-block of dimension 44, where $f$ is the minimum polynomial of an element of order $(q+1)/2$ over $\mathbb{F}_{q}$. (A rational $f$-block refers to an indecomposable $\mathbb{F}_{q}\langle g\rangle$-submodule whose minimum polynomial is $f$.) In particular, the dimension of the centralizer of $g$ in $\mathrm{Sp}_{n}(K)$ is $2+4=6$ (see [56, Theorem 3.1]), and $|C_{G}(g)|=q^{4}|\mathrm{GU}_{1}(q)||\mathrm{O}^{\pm}_{2}(q)|\leq 2q^{4}(q+1)^{2}$ and we see that $|M||C_{G}(g)|^{3}<|G|$.
The other cases are similar. These are $G=\mathrm{SU}_{n}(q)$ and $M$ of type $\mathrm{GU}_{m}(q)\wr S_{t}$; $G=\Omega^{+}_{n}(q)$ and $M$ of type $\mathrm{O}_{m}^{\pm}(q)\wr S_{t}$ ($q$ odd) or $\mathrm{Sp}_{m}(q)\wr S_{t}$ (qtqt even); $G=\Omega_{n}(q)$ with qnqn odd and $M$ of type $\mathrm{O}_{m}(q)\wr S_{t}$. Except for the case where $M$ is of type $\mathrm{O}_{m}^{\pm}(q)\wr S_{t}$, we can choose $g=x\otimes\cdots\otimes x\otimes y$ where $x$ is semisimple with distinct eigenvalues and $y$ unipotent with one Jordan block, as above. In the exceptional case, we have $m\geq 4$ and we choose $y$ unipotent with two Jordan blocks of size $m/2$. We have that all generalized Jordan blocks of $g$ over $K$ have size $m/2$, and each such block occurs with multiplicity at most $2m^{t-2}$, so the dimension of the centralizer of $g$ in $\mathrm{GL}_{n}(K)$ (resp. $\mathrm{O}_{n}(K)$) is at most $2m^{2t-2}$ (resp. at most $m^{2t-2}$, see [56, Theorem 3.1]), and the conclusion follows. ∎
$8.$ Class $\mathcal{C}_{8}$
@@
The cases $G=\mathrm{SL}_{n}(q)$ and $M$ of type $\mathrm{O}^{-}_{n}(q)$ or $\mathrm{Sp}_{n}(q)$ ar similar; in both cases we choose $g$ of order $q^{n/2}+1$. Finally, for $G=\mathrm{Sp}_{n}(q)$ and $M=\mathrm{O}^{\varepsilon}_{n}(q)$ with $\varepsilon\in\{+,-\}$ and $q$ even, letting $g\in M$ be an element without eigenvalue 11, we have $\mathrm{fix}(g,G/M)=1$. ∎
@@ 8.9. Class $\mathcal{N}$
Assume first $G=\mathrm{P}\Omega^{+}_{8}(q)$; we go through the possibilities given in [7, Table 8.50], which is taken from [36]. For convenience, we work in $G=\Omega^{+}_{8}(q)$, and put d=(2,q1)d=(2,q-1). Using the notation as in the table, we are only concerned with novelties $\mathrm{N_{1}}$ and $\mathrm{N_{4}}$. If $M$ is parabolic, we use Lemma 6.3. Assume now $M\cong d\times G_{2}(q)$. We choose $g\in\mathrm{SL}_{3}(q)<M$ of order a ppd of $q^{3}-1$. Then $g$ preserves a nondegenerate 66-space $W$ of plus type (acting irreducibly on complementary maximal totally singular subspaces), so $|C_{G}(g)|\leq(q^{3}-1)(q-1)$. Moreover $|C_{M}(g)|=d(q^{3}-1)/(q-1)$ and $g^{G}\cap M=g^{M}$ (note all cyclic subgroups of $G_{2}(q)$ of order $|g|$ are conjugate). Therefore
@@
Assume next $M=(\Omega^{+}_{2}(q)\times\frac{1}{d}\mathrm{GL}_{3}(q)).[2d]$ with $q\geq 3$; then $M$ is the stabilizer of $\{A,B\}$ where AA and $B$ are totally singular 33-spaces with trivial intersection and $A\oplus B$ nondegenerate; see [36, proof of Proposition 3.2.3]. The subgroup $H=\Omega^{+}_{2}(q)\times\frac{1}{d}\mathrm{GL}_{3}(q)$ preserves AA and $B$. We let $g\in\mathrm{SL}_{3}(q)<H$ be of order a ppd of $q^{3}-1$, acting trivially on $(A\oplus B)^{\perp}$. Note that if $g$ lies in a conjugate of $M$, then it lies in the corresponding conjugate of $H$; but $g$ fixes precisely two totally singular 33-spaces and so $\mathrm{fix}(g,G/M)=1$.
Next let $M=(\Omega^{-}_{2}(q)\times\frac{1}{d}\mathrm{GU}_{3}(q)).[2d]$ with $q\geq 3$. The subgroup $H=\Omega^{-}_{2}(q)\times\frac{1}{d}\mathrm{GU}_{3}(q)$ is the stabilizer in $\mathrm{GU}_{4}(q^{2})\cap G$ of a nondegenerate 11-space for the unitary geometry (see [36, proof of Proposition 3.2.2]). We let $g\in H$ be of order a ppd of $q^{6}-1$ if $q\neq 2$, and of order 99 if $q=2$. Similarly to the previous case we have $\mathrm{fix}(g,G/M)=1$.
Assume finally $M=(D_{2(q^{2}+1)/d})^{2}[2d].S_{2}$. Then $M$ is the normalizer of a Sylow $r$-subgroup of $G$, where $r$ is an odd prime divisor of $q^{2}+1$; we have that $M$ stabilizes a decomposition $V=V_{1}\perp V_{2}$ into 44-spaces of minus type (see [36, Proof of Proposition 3.3.1]). Let $g\in M$ be of order $(q^{2}+1)/d$, acting irreducibly on $V_{1}$ and centralizing $V_{2}$. Then $\mathrm{fix}(g,G/M)$ is equal to the number of Sylow $r$-subgroups of $\Omega(V_{2})\cong\mathrm{PSL}_{2}(q^{2})$, which is $q^{2}(q^{2}-1)/2$, and
@@
Assume now $G=\mathrm{Sp}_{4}(q)$ with $q$ even; the classes in $\mathcal{N}$ can be found in [2, Section 14] (see also [7]). If $M$ is parabolic, then we use Lemma 6.3. If $M=(C_{q-1})^{2}\colon D_{8}$, then $M$ is the normalizer of a nondegenerate 22-space of plus type in $\mathrm{SO}^{+}_{4}(q)<G$. We let $g\in C_{q-1}^{2}<M$ be of order q1q-1 with eigenvalues $\lambda^{\pm 1},\lambda^{\pm 2}$ on the natural module $V$; then $\mathrm{fix}(g,G/M)=1$, since $g$ fixes only two nondegenerate 22-spaces. Assume next $M=(C_{q+1})^{2}\colon D_{8}$, so $M$ is the normalizer of a nondegenerate 22-space of minus type in $\mathrm{SO}^{+}_{4}(q)<G$. We let $g\in M$ be of order $q+1$, with distinct eigenvalues on $V$, and similarly to the previous case we have $\mathrm{fix}(g,G/M)=1$. Assume finally $M=C_{q^{2}+1}\colon 4$, so $M$ is the normalizer of a subgroup $C_{q^{2}+1}$. Then we let $g\in M$ be of order $q^{2}+1$, and we have $\mathrm{fix}(g,G/M)=1$ as in the proof of Lemma 8.6. The proof is now complete. ∎
@@ 9. Classical groups: class $\mathcal{S}$
In this section we complete the proof of Theorem 3.1. The only remaining case is when $G$ is classical and $M$ is in the class $\mathcal{S}$ of subgroups. Throughout this section, $V=\mathbb{F}_{q}^{n}$ denotes the natural module for $G$, where $q$ is a power of the prime pp. (This is a change of notation from Section 8, where the natural module is $\mathbb{F}_{q^{u}}^{n}$ – so in this section, the unitary case is $G=\mathrm{PSU}_{n}(q^{1/2})$.) Class $\mathcal{S}$ consist of the members $M$ of $\mathcal{A}(G)$ that do not belong to $\mathcal{C}_{1},\ldots,\mathcal{C}_{8},\mathcal{N}$. It was proved by Aschbacher [2] (see also [49]) that these subgroups $M$ are almost simple. Moreover, if $L$ denotes the quasisimple cover of $S=\text{Soc}(M)$ acting faithfully on $V$, then $V$ is an absolutely irreducible $\mathbb{F}_{q}L$-module, which cannot be realized over a proper subfield of $\mathbb{F}_{q}$, and finally, if $G=\mathrm{PSL}_{n}(q)$ then $L$ fixes no nondegenerate unitary or bilinear form on $V$.
@@ 9.2. Some preliminary lemmas
Let $a$ be the dimension of the 11-eigenspace of $g$, let $d$ be the dimension of the centralizer of $g$ in $\mathrm{GL}_{n}(K)$ and let $E_{2}$ be the number of distinct irreducible factors of degree at least 22 of the characteristic polynomial of $g$. By Lemmas 7.10 and 7.11 we have $|C_{G}(g)|\leq q^{d/2+a/2+E_{2}}$. Next, note that $E_{2}$ is at most half the number of nontrivial distinct eigenvalues of $g$. Assume first (9.5), so by Lemma 7.7 (with $A=cn$) we have $d+a+2E_{2}\leq B+n^{2}/B$. Therefore, by Lemma 7.9 we get
@@ 9.3. The proof
We now fix some notation. Recall that $G$ is a simple classical group with natural module $V=\mathbb{F}_{q}^{n}$, where $q$ a power of the prime pp. We set $K=\overline{\mathbb{F}_{p}}$. We have that $M\in\mathcal{A}$ is in class $\mathcal{S}$, so $M$ is almost simple; we denote by $S$ the socle of $M$. We will find it convenient, in the proofs, to replace $G$ by the quasisimple cover acting faithfully on $V$, namely $\mathrm{SL}_{n}(q)$, $\mathrm{SU}_{n}(q^{1/2})$ or $\mathrm{Sp}_{n}(q)$ or $\Omega^{\varepsilon}_{n}(q)$. We will still denote by $M$ the preimage in $\mathrm{SL}_{n}(q)$, while we will denote by $L\leq\mathrm{SL}_{n}(q)$ a quasisimple cover of $S$ with $S=L/Z(L)$. For our choice of the element $g\in L$, we will keep throughout the following notation:
@@
Note that $M$ preserves a nondegenerate quadratic or bilinear form on $V$. Since $M\in\mathcal{A}$, it follows that $G$ is orthogonal or symplectic over $\mathbb{F}_{p}$; see for example [39, p. 187] for the precise embeddings. We ignore the cases $G=\mathrm{Sp}_{4}(2)^{\prime}\cong A_{6}$, $G=\Omega^{-}_{4}(2)\cong\mathrm{SL}_{2}(4)$, $G=\Omega_{5}(2)\cong\mathrm{PSp}_{4}(2)$, andqw $G=\Omega^{\pm}_{6}(2)\cong\mathrm{PSL}^{\pm}_{4}(2)$, which have been handled in Lemma 8.1. Assume first $n=4$, so by [7] we have $G=\Omega^{-}_{4}(q)\cong\mathrm{PSL}_{2}(q^{2})$ and $M=A_{5}$ with $q=p\geq 7$. Let $g\in M$ be of order 33, so $|C_{G}(g)|=(q^{2}-1)/2$ and by Lemma 2.1 we have $\mathrm{fix}(g,G/M)=(q^{2}-1)/6<|G:M|^{1/3}$. Assume now $n=5$, so $G=\Omega_{5}(q)\cong\mathrm{PSp}_{4}(q)$ and either $A_{6}\leq M\leq S_{6}$ with $q=p\geq 5$ or $M=A_{7}$ with $q=p=7$. If $M=A_{7}$ we let $g\in M$ be of order 77, so $g$ is regular unipotent and $|C_{G}(g)|=q^{2}$ and so $\mathrm{fix}(g,G/M)\leq 2q^{2}/7<|G:M|^{1/3}$. If $M\leq S_{6}$ then we let $g\in M$ be of order 55, so $g$ is regular in $\mathrm{SL}_{n}(q)$ and $|C_{G}(g)|\leq(q+1)^{2}$ and $\mathrm{fix}(g,G/M)\leq 2(q+1)^{2}/5<|G:M|^{1/3}$. Assume then $n=6$, so $G=\Omega^{\pm}_{6}(q)\cong e.\mathrm{PSL}^{\pm}_{4}(q)$ and $M=e\times A_{7}$ with $q=p\geq 11$ (here $e\in\{1,2\}$). Let $g\in M$ be of order 77, so $|C_{G}(g)|\leq e(q+1)^{3}/2$; moreover $g^{G}$ splits into at most two $M$-classes and so by Lemma 2.1 we have $\mathrm{fix}(g,G/M)\leq(q+1)^{3}/7<|G:M|^{1/3}$. The case $n\geq 7$ is analogous. Specifically, let $g\in A_{m}$ be an $(m-\delta)$-cycle, where $\delta=1$ if $m$ is even and $\delta=0$ if $m$ is odd. In both cases, $g$ is regular in $\mathrm{SL}_{n}(q)$, and $g^{G}$ splits into at most two $M$-classes, so $\mathrm{fix}(g,G/M)\leq 2|C_{G}(g)|/(m-\delta)$, which is easily seen to be $|G|^{1/3}/((2,q-1)\cdot m!)^{1/3}\leq|G:M|^{1/3}$. For example, for $n=8$ and $G=\mathrm{Sp}_{8}(2)$ and $A_{10}\leq M$, $g$ acts on $\mathbb{F}_{2}^{8}$ with irreducible submodules of dimensions 22 and 66, so $|C_{G}(g)|=(2^{3}+1)(2+1)=27$ and
@@
If $n$ satisfies the above inequalities, then by Lemmas 9.1, 9.2, 9.3 and 9.4 we have $|C_{G}(g)|<|G:M|^{1/3}$, as desired. Assume then that $n$ is smaller than the above values. Then, by [32], we see that $m\leq 14$. If $m\leq 13$, then the modular character tables of $L$ are available in [8] and the result follows. If $m=14$, then the modular character tables are available unless $L=2.A_{14}$. By [32], we see that in this case the only option is $n=32$, $p=7$ and $G=\mathrm{Sp}_{32}(q)$. By Lemma 7.17 we have that $L=A_{14}$ is generated by two conjugates of an element $g$ of order $11$. We deduce that either $a\leq 12$ and each nontrivial eigenspace has dimension at most 22, or $a\leq 2$ and each nontrivial eigenspace has dimension at most 33; it follows that $d\leq 184$, which implies $|C_{G}(g)|<|G:M|^{1/3}$. ∎
@@
Assume first $S$ is in $\mathcal{H}_{1}$. Suppose $S$ is in $\mathcal{H}_{1}$(1); then $L=S$. If $S$ is sporadic, let $g=x_{1}$ be as in Table 7.2. If $S$ is of Lie type and $S\neq\mathrm{PSL}^{\varepsilon}_{m}(r)$, let $g=x_{1}$ be as in Table 7.1. If $S=\mathrm{PSL}^{\varepsilon}_{m}(r)$, set $u=1$ if $\varepsilon=+$ and $u=2$ if $\varepsilon=$-, and let $g$ be an element of order $\Phi^{*}_{um}(r)$ if $m$ is odd, and of order $\Phi^{*}_{u(m-1)}(r)$ if $m$ is even. Then, in all cases, [27, Theorem 1.1, Props 3.4–3.13, Prop 4.5] shows that $L$ is generated by three conjugates of $g$ whose product is 11. It follows from Lemma 7.3 that $a\leq n/3$, as desired.